OpenStudy (anonymous):

solve the system xy^3=10^14 x^2/y=10^7

OpenStudy (anonymous):

Solve one equation for a certain variable (in this case I'd be inclined to solve the first one for x), plug that value of x into your second equation, and solve for y. Then, with that value of y, substitute it into any of your equations, and you'll get your value of x.

OpenStudy (anonymous):

y=x^2/10^7 put in eq 1 x(x^2/10^7)^3=10^14 x^7=10^14*10^21=10^35=(10^5)^7 x=10^5 y=10^3